OpenStudy (anonymous):

y= x^3ln (2x+1) ; x=0

OpenStudy (anonymous):

y= x^3 * ln (2x+1) y= (0)^3 *ln (2(0)+1) y= (0)* ln (2(0)+1) y=0

OpenStudy (anonymous):

but 0^0 is not determined !!

OpenStudy (anonymous):

we should probably talk about the limit when x->0 and not x=0 !

OpenStudy (anonymous):

your probabaly right, its limits and I didn't realize it